OpenStudy (anonymous):

Integrate using partial fractions:

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\frac{ 1 }{ \sqrt{x}-\sqrt[3]{x}} dx\]

zepdrix (zepdrix):

hmmmmmmmmmmm....

OpenStudy (abb0t):

Wait a minute, I've seen this question before. Lol

OpenStudy (abb0t):

Must you use PFD?

OpenStudy (anonymous):

Yup, if you have seen this question, you might've seen it in the textbook "Calculus of a Single Variable".

OpenStudy (loser66):

let u = x ^ (1/6) , x = u^6 , dx = 6u^5 du so, x ^(1/2) = u^3 and x^(1/3) = u ^2

OpenStudy (loser66):